Suppose that after 10 years of service, 40% of computers have problems with motherboards (MB), 30% have problems with hard drive (HD), and 15% have problems with both MB and HD. What is the probability that a 10-year old computer still has fully functioning MB and HD?
Answer by jrfrunner(365)

let MB=prob with motherboards
let HD problems with hard drive
P(problems with MB or HD)=P(MB U HD)=P(MB)+P(HD)-P(MB n HD)=0.4+.3-.15=.7-.15=.55
--
This means that P(no MB or HD)= 1-0.55=0.45